# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 9 Related effects
In this section, we describe several kinds of anomalous effects and concepts related with the Mpemba Effect. In particular, we show how different strategies and protocols can affect the relaxation rhythm of a given system. In Sec. 9.1, we discuss the Kovacs effect, where internal states affect the dynamics when the protocol is generalized to more than one temperature quench. The Kovacs effect appears when a naive strategy with two quenches is considered, namely, for example to try to reach faster a given cold temperature going first to a much colder one. A first quench is done to a very cold temperature and then suddenly when the system is at the desired temperature a second quench is applied at that temperature. Next, in Sec. 9.2, we consider various types of asymmetries between heating and cooling quenches, that is, heating can be faster than cooling and vice versa. Specifically, we discuss different results and ways to characterize the existence of an asymmetry between cooling and heating in systems far from equilibrium. In Sec. 9.3, we describe how changes in the ordering of eigenvalues (and their associated eigenvectors) impact the Mpemba effect. And finally, in Sec. 9.4, we show some other mechanisms that can lead to speed-ups in the relaxation to thermal equilibrium. The first consists of to gain leverage using the non-monoticity of equilibrium magnitudes and a weak version of the master equation. Using that approach asymmetry between heating and cooling, preaheating strategy (analogous to precooling strategy [65]) and Mpemba effect are reproduced. The second addresses how to use the evolution of the internal structure of magnetic materials to reach faster a target temperature. In this case, it is shown that the size of internal magnetic domains affects to the velocity of the relaxation after a quench.
@@ 9.1 Kovacs effect and nonmonotonic temporal relaxation
Our discussion of the Mpemba effect has considered situations where a single sudden temperature step, or quench, is applied to a system. Although the relaxation to equilibrium is monotonic in time for a given initial temperature (Fig. 47a), the overall time to relax to equilibrium is—surprisingly—not necessarily a monotonic function of the temperature difference between initial and equilibrium temperatures (Fig. 32). A natural generalization is to ask whether more complicated temperature protocols also lead to new effects. Indeed, at roughly the same time that Mpemba and Osborne were studying water, A. J. Kovacs [247] published an influential study of polymer relaxation showing that a temperature protocol with two temperature steps could lead to nonmonotonic relaxation of the specific volume as a function of time (Fig. 47b). In the original work, the response was a positive overshoot that has become known as the “Kovacs hump”; other systems show a negative overshoot (Fig. 47b).
@@
In glassy materials such as glucose or polyvinyl acetate, Kovacs considered a protocol where the system is prepared in equilibrium at temperature $T_{\mathrm{hot}}$, quenched for a short time at a temperature $T_{\mathrm{cold}}$ and then quenched for a long time at temperature $T_{\mathrm{warm}}$. The switch from cold to warm temperatures is done when the measured response equals the equilibrium value for $T_{\mathrm{warm}}$.222 In the original Kovacs protocol, the time spent at $T_{\mathrm{cold}}$ was only approximately the time for the response to equal its value at $T_{\mathrm{warm}}$. The observed phenomena are qualitatively similar but depend on the exact dwell time. Modern studies simplify the analysis and interpretation by choosing the dwell time as the time for the response to cross the equilibrium value at $T_{\rm warm}$. Naively, if the system is close enough to equilibrium to be described by a single relaxation rate and if the switch to $T_{\mathrm{warm}}$ occurs when the time-dependent response crosses the equilibrium value at $T_{\mathrm{warm}}$, it should stay at that value. However, in systems with multiple degrees of freedom, the previous history—here, the first step—can leave other variables out of equilibrium. The result is an over- or undershoot of the response before a final relaxation to equilibrium. Measuring the response via the fractional volume, Kovacs found overshoots (a single oscillation) in the relaxation for certain combinations of temperatures and times. With colleagues, he proposed an explanation in terms of the multiple rates of relaxation encountered in these highly viscous, glassy systems [248].
Figure 47: Illustration of Kovacs effect. (a) Single-step protocol (Mpemba): A sudden temperature quench $T_{\mathrm{hot}}\searrow T_{\mathrm{warm}}$ (black lines) leads to a monotonic relaxation response (red curve). (b) Double-step protocol (Kovacs): A protocol of $T_{\mathrm{hot}}\searrow T_{\mathrm{cold}}\nearrow T_{\mathrm{warm}}$ leads to a nonmonotonic response (red curve) with a “Kovacs hump” (pink shaded region). The figure is inspired by [249].
@@
Since the original work of Kovacs, investigations have developed in ways reminiscent of the development of our understanding of the Mpemba effect and established two important conclusions: first, nonmonotonic relaxation is a generic behavior of complex systems; second, even very simple systems can nonetheless display Kovacs effects. The overall picture arising from these studies emphasizes the important role of memory and multiple rates.
@@
Observations in other systems. The original experiments by Kovacs were in polyvinyl acetate, a glassy polymer, and in glycerol, a small molecule that shows highly viscous, glassy behavior at low temperatures. It was then natural to look for the Kovacs effect in other complex materials and systems. For example, Mossa and Sciortino found and studied the Kovacs effect in a molecular-dynamics model of orthoterphenyl [250], a small molecule whose relaxation is qualitatively similar to that of glycerol. Later, Chen et al. applied the Kovacs protocol to a colloidal glass of nanospheres with temperature-dependent, attractive interactions [95]. Using X-ray photon correlation spectroscopy (XPCS), they found nonmonotonic responses in quantities such as the elastic and viscous response functions.
@@
The Kovacs effect has also been observed in still other settings. Josserand et al. observed overshoot in the relaxation of the level of a granular fluid (glass beads in a vertically vibrated cylinder) after a Kovacs protocol in the vertical acceleration [256]. Kürsten et al. studied a 2D model of interacting particles inspired by studies of bacterial swarms [257]. They applied a temperature-step protocol and monitored the response of the orientational order of the particles. Finally, Prados and Trizac showed the existence of the Kovacs effect and its inverse in granular gases [258, 259] and Lasanta et al. studied numerically a granular gas of rough spheres and found not only unusually large, order-one Kovacs response but also responses with two and even three extrema during the relaxation [260]. Interestingly, Mompó et al. [142] showed in a gas of viscoelastic granular particles a unified picture of the Mpemba and Kovacs effects based on the evolution of kurtosis see Sec. 4.
@@
A 2021 experiment by Militaru et al. observed experimentally a Kovacs effect (also in mean energy response) for the relaxation of a nanoparticle trapped by optical tweezers in air [249]. In contrast to all the previous observations of the Kovacs effect, there was just a single particle with a single relaxing mode. The pressure (near ambient) created a nearly critical damping level, where the natural oscillation frequency and relaxation rate were comparable. In the experiment, an effective temperature was manipulated by subjecting the bead to random fluctuating electrostatic forces [263]. The ensemble-averaged potential energy relaxation response showed a nonmonotonic response for nearly critical damping levels (shaded region in Fig. 47b). Indeed, at the time of the switch to $T_{\mathrm{warm}}$, the average potential energy equaled its equilibrium value, but the average kinetic energy did not. The hump reflects the subsequent equilibration of both kinetic and potential energies. Interestingly, the effect vanishes when the vibration and damping rates are mismatched (limits of weak damping or strong overdamping).
Given the complexity of the materials where the Kovacs effect has been typically studied, it is not surprising that there has been a wide range of theoretical studies, e.g., Refs. [248, 264, 153, 265]. A common feature is the role of multiple, comparable rates and of memory [136]. Indeed, the Kovacs effect has been considered as a simple example of the importance of memory effects (initial conditions and history) as a means of creating and engineering matter with desirable new properties [266]. Now that we understand that the Kovacs effect can arise in systems with a few comparable rates, it is perhaps not surprising that it is so widespread in Nature and in theoretical models.
@@ 9.2 Asymmetric heating and cooling
Although the above discussion suggests that heating and cooling times are identical for initial temperatures that are infinitesimally different from the bath, relaxation times might differ when the initial states are far from equilibrium. Mathematically, the function $a_{2}(T_{0},T_{b})$ need not be odd about $T_{b}$, even though it must vanish at $T_{0}=T_{b}$. Following this line of thought, Lapolla and Godec used a nonequilibrium free energy difference $F(t)$ to quantify the distance between the initial state and final equilibrium. In particular, if we consider three temperatures $T_{c}$, $T_{w}$, and $T_{h}$ (cold, warm, hot) and let $i\in\{h,c\}$, then [270, Sec. 5.2]
@@
measures the non-equilibrium free energy difference between the relaxing hot or cold state $p_{i}(x,t)$ and the (warm) equilibrium bath state characterized by $\pi(x,T_{w})$, which is related to the Kullback-Leibler divergence $D_{\rm KL}$ and defines a good way to compare two different states of the system by means of its probability distributions [271].
@@
Figure 48: Heating is faster than cooling for equidistant quenches. (a) Schematic of experimental setup: A bead in an optical trap with harmonic potential and stochastic electrical forces whose strength quantifies the effective temperature of the bead. (b) Evolution of Gaussian position distributions after quenches from $T_{h}$ or $T_{c}$ to $T_{w}$ (forward-protocol) or vice versa (backward protocol). Thicker lines denote faster relaxation pathways. (c) Schematic of forward and reverse protocols. These figures are reproduces from [273].
Figure 49: Heating is faster than cooling for equidistant quenches. Red arrows represent heating whereas blue ones cooling. Solid lines correspond to the theoretical predictions. Solid and dashed arrows stand for the forward and backward protocols, respectively, and thicker lines indicate the fastest evolution. (a) and (b) show time evolution of the generalized excess free energy for equidistant quenches, (a) are results for the forward protocol and (b) for the backward protocol. Insets represent the initial value of the equidistant relative entropy as a function of the temperature. Figures are adapted from [273].
@@
A natural question to ask is how universal this asymmetry is. Already, Lapolla and Godec noted counterexamples when the potential had metastable states where the relaxation required thermal hops over barriers [268]. Then Meibohm found counterexamples for single-well potentials with non-quadratic forms $U(x)=|x|^{\alpha}$, for $\alpha\gtrsim 3$ [275]. A small additional quadratic contribution to the potential does not alter the result, meaning that there can be counterexamples for wells with strong-enough anharmonic shapes. Van Vu and Hasegawa examined relaxation in classical discrete-state systems [276]. They confirmed the asymmetry for two-state systems but found counterexamples in systems with more than two states. In particular, they found and demonstrated that in more than two-level quantum systems, for a three-temperature protocol with a warm intermediate temperature and for some fitting parameters, the asymmetry can be reversed, and the cooling process can be faster than the heating one. Even for degenerate two-state systems, the asymmetry could be reversed beyond a critical threshold for the energy gap. Also, Manikandan showed that in discrete-state quantum systems, asymmetries are not generic and can have either sign, depending on various system parameters [277].
Figure 50: Thermal kinematics of heating and cooling between any pair of temperatures. Red circles denote heating, blue squares cooling. Solid lines are theoretical predictions. (a) Instantaneous statistical velocity as a function of time. (b) Degree of completion as a function of time. Figure are adapted from [273].
@@
Figure 51: Heating is faster than cooling in a quantum Brownian particle. (a) Simulation of the evolution of the fidelity with respect to the thermal state $T_{w}$ for equidistant initial states. (b) Degree of completion computed for the heating and cooling processes in the two-temperature protocol. For both panels the $x$ axes represent a rescaled time with respect the total time the system takes to reach the final state. (c) Eigenvalues of the operator $\mathcal{L}$ from  Eq. (60). (d) First eigenvalues of (c); marker size is proportional to the overlap with the thermal state at the opposite temperature. Figure are reproduced from [279].

### v3

@@ 9 Related effects
In this section, we describe several kinds of anomalous effects and concepts related with the Mpemba Effect. In particular, we show how different strategies and protocols can affect the relaxation rhythm of a given system. In Sec. 9.1, we discuss the Kovacs effect, where internal states affect the dynamics when the protocol is generalized to more than one temperature quench. The Kovacs effect appears when a naive strategy with two quenches is considered, for example to try to reach faster a given cold temperature by going first to a much colder one. A first quench is to a very cold temperature. Then, suddenly, when the system is at the desired temperature, a second quench is applied at that temperature. In Sec. 9.2, we consider various types of asymmetries between heating and cooling quenches, that is, heating can be faster than cooling and vice versa. Specifically, we discuss different results and ways to characterize the existence of an asymmetry between cooling and heating in systems far from equilibrium. In Sec. 9.3, we describe how changes in the ordering of eigenvalues (and their associated eigenvectors) impact the Mpemba effect. And finally, in Sec. 9.4, we show some other mechanisms that can lead to speedups in the relaxation to thermal equilibrium. The first is to gain leverage using the non-monotonicity of equilibrium magnitudes and a weak version of the master equation. Using that approach, asymmetry between heating and cooling, preheating strategy (analogous to the precooling strategy introduced in [65]) and the Mpemba effect are all reproduced. The second addresses how to use the evolution of the internal structure of magnetic materials to reach faster a target temperature. In this case, it is shown that the size of internal magnetic domains affects to the velocity of the relaxation after a quench.
@@ 9.1 Kovacs effect and nonmonotonic temporal relaxation
Our discussion of the Mpemba effect has considered situations where a single sudden temperature step, or quench, is applied to a system. Although the relaxation to equilibrium is monotonic in time for a given initial temperature (Fig. 47a), the overall time to relax to equilibrium is—surprisingly—not necessarily a monotonic function of the temperature difference between initial and equilibrium temperatures; see Fig. 32. A natural generalization is to ask whether more complicated temperature protocols also lead to new effects. Indeed, at roughly the same time that Mpemba and Osborne were studying water, A. J. Kovacs [258] published an influential study of polymer relaxation showing that a temperature protocol with two temperature steps could lead to nonmonotonic relaxation of the specific volume as a function of time; see Fig. 47b. In the original work, the response led to a positive overshoot that has become known as the “Kovacs hump.” Later, Prados and Trizac [259] showed in granular gases that there also can exist an “anomalous” downward Kovacs hump; see Fig. 47a for a sketch of both types of Kovacs response, and see Fig. 47b for an anomalous hump observed in an optically trapped colloidal-particle experiment).
@@
In glassy materials such as glucose or polyvinyl acetate, Kovacs considered a protocol where the system is prepared in equilibrium at temperature $T_{\mathrm{hot}}$, quenched for a short time at a temperature $T_{\mathrm{cold}}$ and then quenched for a long time at temperature $T_{\mathrm{warm}}$. The switch from cold to warm temperatures is done when the measured response equals the equilibrium value for $T_{\mathrm{warm}}$.222 In the original Kovacs protocol, the time spent at $T_{\mathrm{cold}}$ was only approximately the time for the response to equal its value at $T_{\mathrm{warm}}$. The observed phenomena are qualitatively similar but depend on the exact dwell time. Modern studies simplify the analysis and interpretation by choosing the dwell time as the time for the response to cross the equilibrium value at $T_{\rm warm}$. Naively, if the system is close enough to equilibrium to be described by a single relaxation rate and if the switch to $T_{\mathrm{warm}}$ occurs when the time-dependent response crosses the equilibrium value at $T_{\mathrm{warm}}$, it should stay at that value. However, in systems with multiple degrees of freedom, the previous history—here, the first step—can leave other variables out of equilibrium. The result is an over- or undershoot of the response before a final relaxation to equilibrium. Measuring the response via the fractional volume, Kovacs found overshoots (a single oscillation) in the relaxation for certain combinations of temperatures and times. With colleagues, he proposed an explanation in terms of the multiple rates of relaxation encountered in these highly viscous, glassy systems [260]. A general explanation using the master-equation formalism [86] was given in [161, 261], where it was shown that the normal Kovacs effect is produced in the linear-response regime.
Figure 47: Illustration of Kovacs effect. (a) Sketch of the normal and anomalous Kovacs response produced in a driven granular gas; see Sec. 4. Top: Driving protocol has three forcing levels, $\xi_{0}$, $\xi_{1}$, and $\xi_{2}$. Bottom: Kinetic temperature time evolution. Initially, the system is at the hot temperature $T_{s}(\xi_{0})$, in a nonequilibrium steady state resulting from the driving $\xi_{0}$. Then, the driving level is suddenly highly decreased to $\xi_{1}$, and the system would asymptotically reach the cold temperature $T_{s}(\xi_{1})$. However, at $t=t_{w}$, the granular temperature is measured and the driving modified to $\xi$. The new value would lead to a warm temperature, $T_{s}(\xi)=T(t_{w})$. The question mark illustrates the two possible Kovacs hump, a positive hump with a maximum (normal behavior, solid line) and a negative anomalous hump (dot-dashed line). (b) Experimental observation of the Kovacs effect in an optically trapped particle. Top: Temperature protocol of the effective bath during the cooling protocol from $T_{H}$ to $T_{C}$ (dashed light blue line), cooling protocol from $T_{H}$ to $T_{W}$ (dash-dotted silver line), and Kovacs protocol (red solid line). Bottom panel: Time evolution of average potential energy during the protocols described in (a). Discrete symbols (light blue triangles, silver diamonds, and red circles) are used for experimental data, whereas lines are used for theory curves. The horizontal dotted line shows the equilibrium value at the warm temperature. The vertical dotted line marks the time $t_{w}$ at which the temperature is changed to $T_{w}$. The plot shows an anomalous (downward) “Kovacs hump.” Source: Reprinted with permission from [259] and [262].
@@
Since the original work of Kovacs, investigations have developed in ways reminiscent of the development of our understanding of the Mpemba effect and have established two important conclusions: first, nonmonotonic relaxation is a generic behavior of complex systems; second, even very simple systems can nonetheless display Kovacs effects. The overall picture arising from these studies emphasizes the important role of memory and multiple rates.
@@
Observations in other systems. The original experiments by Kovacs were in polyvinyl acetate, a glassy polymer, and in glycerol, a small molecule that shows highly viscous, glassy behavior at low temperatures. It was then natural to look for the Kovacs effect in other complex materials and systems. For example, Mossa and Sciortino found and studied the Kovacs effect in a molecular-dynamics model of orthoterphenyl [263], a small molecule whose relaxation is qualitatively similar to that of glycerol, and Berthier and Bouchaud found the effect numerically in an Ising spin glass [264]. Later, Chen et al. applied the Kovacs protocol to a colloidal glass of nanospheres with temperature-dependent, attractive interactions [95]. Using X-ray photon correlation spectroscopy (XPCS), they found nonmonotonic responses in quantities such as the elastic and viscous response functions.
@@
The Kovacs effect has also been observed in still other settings. Josserand et al. observed overshoot in the relaxation of the level of a granular fluid (glass beads in a vertically vibrated cylinder) after a Kovacs protocol in the vertical acceleration [270]. Prados and Trizac showed for the first time the existence of the normal and anomalous Kovacs effect in granular gases [259, 271]; Lasanta et al. studied numerically a granular gas of rough spheres, finding not only unusually large, order-one Kovacs response but also responses with two and even three extrema during the relaxation [272]. Motivated by the results in granular gases, Kürsten et al. studied a 2D model of interacting particles inspired by studies of bacterial swarms [273]. They applied a temperature-step protocol and monitored the response of the orientational order of the particles. Interestingly, Mompó et al. [149] showed in a gas of viscoelastic granular particles a unified picture of the Mpemba and Kovacs effects based on the evolution of kurtosis; see Sec. 4.
@@
A 2021 experiment by Militaru et al. observed experimentally an anomalous Kovacs effect, also in mean energy response, for the relaxation of a nanoparticle trapped by optical tweezers in air [262]; see Fig. 47b). In contrast to previous observations of the Kovacs effect, there was just a single particle with a single relaxing mode. The pressure (near ambient) created a nearly critical damping level, where the natural oscillation frequency and relaxation rate were comparable. In the experiment, an effective temperature was manipulated by subjecting the bead to random fluctuating electrostatic forces [276]. The ensemble-averaged potential energy relaxation response showed a nonmonotonic response for nearly critical damping levels; shaded region in Fig. 47b. Indeed, at the time of the switch to $T_{\mathrm{warm}}$, the average potential energy equalled its equilibrium value, but the average kinetic energy did not. The hump reflects the subsequent equilibration of both kinetic and potential energies. Interestingly, the effect vanishes when the vibration and damping rates are mismatched, i.e., in the limits of weak damping or strong overdamping.
Given the complexity of the materials where the Kovacs effect has been typically studied, it is not surprising that there have been many theoretical studies, e.g., Refs. [260, 277, 161, 278]. A common feature is the role of multiple, comparable rates and of memory [143]. Indeed, the Kovacs effect has been considered as a simple example of the importance of memory effects (initial conditions and history) as a means of creating and engineering matter with desirable new properties [279]. Now that we understand that the Kovacs effect can arise in systems with a few comparable rates, it is perhaps not surprising that it is so widespread in Nature and in theoretical models.
@@ 9.2 Asymmetric heating and cooling
Although the above discussion suggests that heating and cooling times are identical for initial temperatures that are infinitesimally different from the bath, relaxation times might differ when the initial states are far from equilibrium. Mathematically, the function $a_{2}(T_{0},T_{b})$ need not be odd about $T_{b}$, even though it must vanish at $T_{0}=T_{b}$. Following this line of thought, Lapolla and Godec used a nonequilibrium free energy difference $F(t)$ to quantify the distance between the initial state and final equilibrium. In particular, if we consider three temperatures $T_{c}$, $T_{w}$, and $T_{h}$ (cold, warm, hot) and let $i\in\{h,c\}$, then [283, Sec. 5.2]
@@
measures the nonequilibrium free energy difference between the relaxing hot or cold state $p_{i}(x,t)$ and the (warm) equilibrium bath state characterized by $\pi(x,T_{w})$, which is related to the Kullback-Leibler divergence $D_{\rm KL}$ and defines a good way to compare two different states of the system by means of its probability distributions [284].
@@
Figure 48: Heating is faster than cooling for equidistant quenches. (a) Schematic of experimental setup: A bead in an optical trap with harmonic potential and stochastic electrical forces whose strength quantifies the effective temperature of the bead. (b) Evolution of Gaussian position distributions after quenches from $T_{h}$ or $T_{c}$ to $T_{w}$ (forward-protocol) or vice versa (backward protocol). Thicker lines denote faster relaxation pathways. (c) Schematic of forward and reverse protocols. Source: Reprinted with permission from [286].
Figure 49: Heating is faster than cooling for equidistant quenches. Red arrows represent heating whereas blue ones cooling. Solid lines correspond to the theoretical predictions. Solid and dashed arrows stand for the forward and backward protocols, respectively, and thicker lines indicate the fastest evolution. (a) and (b) show time evolution of the generalized excess free energy for equidistant quenches, (a) are results for the forward protocol and (b) for the backward protocol. Insets represent the initial value of the equidistant relative entropy as a function of the temperature. Source: Adapted with permission from [286].
@@
A natural question to ask is how universal is this asymmetry. Already, Lapolla and Godec noted counterexamples when the potential had metastable states where the relaxation required thermal hops over barriers [281]. Then Meibohm found counterexamples for single-well potentials with non-quadratic forms $U(x)=|x|^{\alpha}$, for $\alpha\gtrsim 3$ [288]. A small additional quadratic contribution to the potential does not alter the result, meaning that there can be counterexamples for wells with strong-enough anharmonic shapes. Van Vu and Hasegawa examined relaxation in classical discrete-state systems [289]. They confirmed the asymmetry for two-state systems but found counterexamples in systems with more than two states. In particular, they found and demonstrated that in more than two-level quantum systems, for a three-temperature protocol with a warm intermediate temperature and for some fitting parameters, the asymmetry can be reversed, and the cooling process can be faster than the heating one. Even for degenerate two-state systems, the asymmetry could be reversed beyond a critical threshold for the energy gap. Also, Manikandan showed that in discrete-state quantum systems, asymmetries are not generic and can have either sign, depending on various system parameters [290].
Figure 50: Thermal kinematics of heating and cooling between any pair of temperatures. Red circles denote heating, blue squares cooling. Solid lines are theoretical predictions. (a) Instantaneous statistical velocity as a function of time. (b) Degree of completion as a function of time. Source: Adapted with permission from [286].
@@
Figure 51: Heating is faster than cooling in a quantum Brownian particle. (a) Simulation of the evolution of the fidelity with respect to the thermal state $T_{w}$ for equidistant initial states. (b) Degree of completion computed for the heating and cooling processes in the two-temperature protocol. For both panels the $x$ axes represent a rescaled time with respect the total time the system takes to reach the final state. (c) Eigenvalues of the operator $\mathcal{L}$ from  Eq. (60). (d) First eigenvalues of (c); marker size is proportional to the overlap with the thermal state at the opposite temperature. Source: Reprinted with permission from [292].